40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 11 14:10:05 2013, Jenne, Update, Environment, MC suspensions moved by this morning's earthquakes EQ.pngMC-SUS-kicked-EQ.png
    Reply  Mon Mar 11 15:52:46 2013, Jenne, Update, IOO, MC spots centered, WFS realigned MCspots_11Mar2013.png
    Reply  Mon Mar 11 17:18:00 2013, Manasa, Update, Environment, Earthquake: Suspensions tripped and MC realigned 
       Reply  Tue Mar 12 00:45:50 2013, Jenne, Update, IOO, MC weird again 
          Reply  Tue Mar 12 11:49:18 2013, Jenne, Update, IOO, MC weird again 
             Reply  Wed Mar 13 02:48:13 2013, Jenne, Update, IOO, MC all better 
Message ID: 8268     Entry time: Mon Mar 11 14:10:05 2013     Reply to this: 8269   8271
Author: Jenne 
Type: Update 
Category: Environment 
Subject: MC suspensions moved by this morning's earthquakes 

None of the suspensions All suspensions were tripped (edited by Manasa; see elog 8271) by this morning's earthquakes, but the MC suspensions are in a different place than they were a day ago.  The big symptom here is that the MCWFS are pulling the mode cleaner slightly out of alignment.  When it first locks, the reflected light is ~0.5, but when the WFS are engaged it goes up to ~0.8.  I'm going to put the MC optics back where they were (according to SUSpit and SUSyaw), and tweak up the MC from there. Probably other optics are affected, but I was going to work on continuing to center the beam on the Yarm optics, so I'll deal with the rest of the IFO in a minute.

Note re: lower plot - the mxstream was down on c1sus and c1ioo, so no fast channels on those computers were recorded for almost a day.  (The plot is one day 4 days long).  I was going to plot the seismic blrms along with the suspension pointing values, but there's no data saved, so there's no point.  Jamie tells me he thinks this spontaneous loss of the mxstream is fixed in the next RCG upgrade, and that we can talk about upgrading the RCG after the LSC meeting, so this data loss is no longer an issue.

EQ.png

MC-SUS-kicked-EQ.png

EDIT:  Plot with 4 days of trend, rather than just 1.  The MC alignment (as measured by MC refl) has been very bad for several days.  I'm going to move the suspensions back to their last good place.  Also, Manasa realigned the MC after the EQ, so I don't actually know where the suspensions got kicked to this morning.

ELOG V3.1.3-